korosh23:

Discrete Math Question: Given P (n , 3) = 3*P (n , 2) What is the value of n ?

korosh23:

Hint: This question uses permutations P(n,r) = n!/(n-r)! so : n!/(n-3)! = 3 * n!/ (n-2)! How do we go from here ?

myininaya:

Hint (n-2)!=(n-2)*(n-3)!

korosh23:

Yes! I got the right answer! @myininaya could you please tell me how did you know (n-2)! = (n-2)(n-3)! ?

moreso:

$$(n-2)! = (n-2)(n-3)... (3)(2)(1) \\ ~ \\ (n-2)!= (n-2 ) \left( (n-3)(n-4)... (3)(2)(1)\right) \\ \, \\ (n-2)!= (n-2 )(n-3)!$$